In 2015 the amount of fixed investments in Russia was 13,9 billion Rubles. In 2016 the amount of fixed investments increased by 0,8 billion Rubles in comparison with 2015. During the period from 2015 to 2017, the average annual growth rate of this indicator was 1,05 billion Rubles. Calculate the amount of fixed investment in 2017 year. (billion Rubles):


1
Expert's answer
2021-07-26T03:49:02-0400

Fixed investment in 2015=13.9b

Fixed investment in 2016=13.9 + 0.8 b = 14.7b

Fixed investment in 2017= y


Average annual growth rate between 2015-2017= 1.05

Which is growth rate of 2016 and 2017 divided by 2.

Growth rate after 2016 will be

(14.7- 13.9)/13.9=0.0575

Growth rate after 2017 will be given as y. Hence (0.0575+y)/2= 1.05

which is 2.0425

To get fixed investment for 2017 we assume it to be m. Hence (m-14.7)/14.7=2.0425

=20.95